Find the area between the curves y^2 - y + x = 0 and y^2 = x + 3.


1
Expert's answer
2021-05-19T12:50:39-0400

Find the area between the curves "y^2-y+x=0" and "y^2=x+3".

"\textbf{Solution:}"

Let "x_{2}(y)=y-y^2", "x_{1}(y)=y^2-3".

Let's draw the graphs of these functions.



We have two points of intersection.

"y-y^2=y^2-3 \Rightarrow 2y^2-y-3=0".

"y_{1}=1.5 \Rightarrow x_{1}=-0.75".

"y_{2}=-1 \Rightarrow x_{2}=-2".

Find the area "S".

"S=\int_{y_2}^{y_1} (x_{2}(y)-x_{1}(y)) dy."


"S=\int_{-1}^{1.5} ((y-y^2)-(y^2-3)) dy=\int_{-1}^{1.5} (-2y^2+y+3) dy."


"S=(-2)\int_{-1}^{1.5} y^2 dy+\int_{-1}^{1.5} y dy+3\int_{-1}^{1.5} dy=" "(-2)\left.\frac{y^3}{3}\right|_{-1}^{1.5}+\left.\frac{y^2}{2}\right|_{-1}^{1.5}+3\left.y\right|_{-1}^{1.5}=\frac{125}{24}".

"\textbf{Answer:}"

"\boxed{S=\frac{125}{24}}"